# Aristotle’s Second-Order Logic

### v11

@@ 7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
which is topic 5 of Chapter 1 listed in appendix 1. By what we have seen we can assume that such an inference could be considered as obtained via Aristotle’s deployment of natural deduction rules for first and second-order quantifiers as well as propositional equivalences.
Although the topics are usually listed in an ad hoc fashion it is clear that they are organized into groups closely knit together by logical equivalence or implication. We hold that second-order logic is required both to formulate the topics and (in its natural deduction form) to explain how Aristotle might have established the inferences and equivalence therein. We might call the working out of these implicit inferences the ’metalogic’ of the Topics. Our objective here is not to undertake an exhaustive formal interpretation of all the logical topics in ASOL. Rather we will go through a selection of key topics mainly from book IV but also from books V and VI which provide evidence for the second-order nature of the metalogic. Let us start with book IV and $\gamma$. We find therein topics which state directly or are stated to be inferred from $\gamma$ having transitive, symmetric and anti-reflexive properties.
@@
as can be verified in appendix 1. The first topic is clearly an ancestor of the Barbara syllogism of the Analytics. Perhaps $\mu$ descends for instance from $\delta$ since the same difference can operate on distinct genera so that $\delta AB$, $A$ is a difference of genus $B$ means that ’some $B$ are $A$’ ( topic (1) further ahead has certainly some connection to the syllogism Darii). Then there is the important tree-like property:
@@
which is the first topic in chapter 2 of book IV. Aristotle refers to $\gamma AB$ also as $A$ participating in $B$ or $A$ receiving the the definition of $B$. Aristotle mentions important axioms in the context of the derivation of refutation or establishing topics. This gives us insight into the ’metalogic’ of the topics. For instance in 122b Aristotle infers a topic using the tree-like property and the topic can itself be represented as inference in sequent form:
@@
There is the defined concept of proximate genus:
@@
This furnishes us with an example of nested heterogenous quantifiers over non-monadic (second-order) predicates.
Another example: every genus must have more than one proximate species (a species of which it is the proximate genus)
@@
Here we have an appearance of equality for monadic predicate terms, which we should add to our arsenal of second-order predicates - for we cannot reduce it to first-order extensional equivalence ($=_{e}$ of section2). We certainly have
@@
but not (as we argued previously) the reciprocal implication. We could tentatively posit that we have a second-order version of Leibniz’s law (see the topics in 152a which suggests particular cases of this law). In particular in book VII 152a31-33 we read (tr. W. A. Pickard):
@@
Aristotle appears to be stating that since $\sim(A=C\,\&\,B=C)\rightarrow\sim(A=B)$ then a valid topic is
$(A=B)\,\&\,(A=C)\rightarrow(B=C)$
The principle above about the plurality of immediate species furnishes us with important example of nested heterogenous quantifiers over non-monadic (second-order) predicates. Note that it follows from the axioms that the proximate genus of a term is unique (modulo the equality relation).
Axioms ($\Gamma$) and (I) are stated as a topics. So too are often the contrapositives of some of the topics considered above .
Our second-order binary predicates are also exclusive of each other. For instance we have the topic $\sigma AB\rightarrow\sim\gamma AB$ but this can be derived from ($\Sigma$) and ($\Gamma$). We also have that $\sim\Sigma AA$. There is a topic which seems to state that if $\gamma AB$ then $B$ cannot be the difference of a genus of $A$:
@@
If $\gamma AB$ then there must be some difference $C$ of $B$, $\delta CB$ which ’is said’ of $A$. Aristotle has a kind of unspecific general kind of ’predication’ wherein it is not obvious whether it is supposed to fall under one of the predicabilia or not. We take this unspecified predication to be just extensional containment $A\subset B\leftrightarrow\forall x(Ax\rightarrow Bx)$. We might express this topic as
@@
From this topic we derive the refutation topic:
@@
We will not analyze here the question of supreme genera and infima species but the order-theoretic properties involving $\gamma$ could be easily expressed in ASOL though other implied properties such as the finiteness of ascending chains of subordinate genera would have to make use of an extension of ASOL such as Zalta’s Object Logic[25].
We have not addressed to the ontological matter of the different categories which feature in the Topics. For our purposes we can just remark that for the order on monadic predicate terms induced by $\gamma$ there can be in general more than one maximal element. And we can ask if it is correct to interpret our first-order individuals as individual substances. There is also a problem that Aristotle would seem to distinguish different kinds of situation for which $Ax$, for which a predicate term $A$ holds of an individual substance $x$, for example, if $A$ is predicated according to essence or as an accident. This would suggest that we need an extra second-order binary predicate with second-order and first-order arguments $\gamma^{i}Ax$ (for which $\gamma^{i}Ax\rightarrow Ax)$ to make such ontological predication distinctions of individuals. A general approach in ASOL the theory of predication in the The Categories might be something as follows (we will not discuss here the question of whether some of the other categories besides relation are in fact also relations: knowledge is a state of a subject but also, it seems, a relation). We postulate nine different categories of predication $\mathcal{P}^{i}Tx$ where we put $\mathcal{P}^{1}Tx\leftrightarrow Tx$ as corresponding to ousia while the others correspond to the other Aristotelian categories. The tenth will correspond to relation and be $\mathcal{P}^{10}R^{2}xy$ or $\mathcal{P}^{10}R^{2}xT$. For consider ’knowledge’ and ’knowledge of colours’. Knowledge corresponds to the state of a subject but also to a relation, and this relation can regard not one or many primary substances but other categories: if a subject $a$ knows the colour ’blue’ then this is not the same as knowing all blue substances.
A central role in the Topics is played be a kind of ’semantic’ term negation operator (opposite operator), which can be seen as operating within the proximate species of a given genus. This must be a second-order function applied (primarily) to monadic predicate terms. Its fundamental properties are
@@
For instance the opposite of the color white is black. A proximate species of a genus may not have an opposite; this situation which we could naturally express as $A^{\bullet}=A$. Noteworthy topics are
@@
In order to formalize the theory of definition we cannot do without a binary function on monadic predicate terms which corresponds grammatically to a modifier in a noun-phrase, for instance an adjective attached to a head noun. We study in more detail the grammatical aspect of definition terms at the end of this section. We denote the function by means of concatenation and impose the axiom which is implicitly assumed by Aristotle:
@@
Since we do not want semantically meaningless constructs like ”green numbers” it makes sense to define a further second-order monadic predicate allowing us to express non non-empty extension (cf. also (NEmp) in section 2):
@@
From this and the axiom for our concatenation operation we can deduce the that immediate species of a genus have disjoint extensions:
@@
It is important to note that equality here is not to be read syntactically but semantically. Thus $D=BC$ is to be read as ’the concept $D$ has as parts concepts $B$ and $C$’, the paradigm being ’rational animal’. We hold to be a fundamental problem whether a term can be said to be equal to its definition (with this notion of equality) and whether Leibniz’s law holds for this notion of equality. It could be that two distinct kinds of equality are required. It is true however that definition though the substitution of a term by its definition features in several topics202020152a-152b is very relevant for term equality. Aristotle announces topics which express the transitivity of $=$ as well as topics recalling Leibniz’s law. In 153a Aristotle seems to say that there is equality between a term and its definition but this is in itself not enough to show that it is a definition. The definiens must satisfy further conditions. The passage 140b32-141a1 is also interesting. Aristotle states that a definition is identical to the term it defines: esti de tauton tô(i) anthrôpô(i) zô(i)on pezon dipoun. However in 143a15-28 where Aristotle appears to admit that either giving the immediate genus and the difference or a non-immediate genus and all corresponding differences (the ’intension’) as both valid forms of definition which leads us to inquire exactly what the nature of the purported ’uniqueness’ of the definition. Does this means that both these forms are to be regarded as the same definition ?. From this last definition we can evidently derive (in ASOL) numerous refutation topics - and these are indeed topics in book VI. As Aristotle says: it is much easier to refute a definition than to establish one.
A standard view would be that if $hA(DC)$ then $C$ must be uniquely determined (this in fact follows from the above axioms). There is however a problem suggested by the passage 148b dealing with ’complex terms’212121this notion of ’complex term’ appears to be primarily grammatical and not logical or semantic as is our concatenation operation, however the precise relation between the two remains illusive as does the interpretation of the passage in question. Are we to understand the crucial non-circularity property of the definiens in semantic or syntactic terms ? But semantically does it even make sense to have a situation in which $A=AB$
@@
It is also interesting how definitions relates to the opposite operator. If we have a definition $DG$ of $A$ what can we say in general about the definition of $A^{\bullet}$ In 153b Aristotle says that there are three cases to consider for the definition for $A^{\bullet}$ depending on whether $G^{\bullet}=G$ or $D^{\bullet}=D$. The possibilities are either $D^{\bullet}G^{\bullet}$, $D^{\bullet}G$ or $DG^{\bullet}$. For example justice is virtue of the soul and its opposite injustice is the vice of the soul, soul having no opposite.
We now discuss Aristotle’s relation terms concerning which section 4 of book IV of the Topics is of particular interest. A natural approach is to represent relation terms by binary predicate terms in ASOL. But then we need a corresponding version of $\gamma$ for such arguments, there is no way to deal uniformly with the predicabilia when predicate terms of different arities are involved. This is also enforced by the topic in book IV section 1 (120b) which requires that species and genus belong to the same category, although Aristotle admits later in section 4 that there might be exceptions (124b) as for the case of ’grammatical’ (monadic) as a species of the genus ’knowledge’(relation). In our case the application of the aforementioned topic (stated also explicitly in (124b)) is simply that there must be two species-genus relations, one $\gamma$ for monadic predicates and another $\gamma_{2}$ for binary predicate variables. In a formalization we will be required to postulate analogous axioms as for $\gamma$, for instance:
@@
where $R,S$ and $T$ are binary predicate variables. The logic of relation terms is quite complex in Aristotle and bound up with Greek syntax in a way which renders interpretation difficult. One thing that is clear is that for a relation term $R$ we have its ’reciprocal’ or ’inverse’ which we can interpreted by a monadic second-order function on binary predicate terms thus $R^{\ast}$. A challenge is to interpret the passage in (124b) about the double ($D$) being the double of the half. This suggests first of all that Aristotle is considering one-to-one relations
$\forall x,x^{\prime},y,y^{\prime}(Rxy\&Rxy^{\prime}\rightarrow x=y^{\prime})\&(Rxy\&Rx^{\prime}y\rightarrow x=x^{\prime})$
@@
It would be interesting to investigate the connection between the theory of relations in the Topics and the interpretation of ekthesis in terms of $\epsilon$ and the generalization of $\iota$ discussed in section 3.
In the passages that follow Aristotle discusses the necessary agreement between the case governed for $\gamma_{2}$. For example if $\gamma_{2}RS$ then if $R$ takes the genitive then so does $S$. More interesting from a logical point of view is the passage mentioning the number of complements, this suggests that Aristotle was conscious of the necessity of the general agreement between the arity of the predicate terms in the species-genus relation, so that we should consider a whole sequence of such relations $\gamma_{1}=\gamma,\gamma_{2},...,\gamma_{i},....$ where $\gamma_{n}$ serves for $n$-ary predicate variables. Aristotle uses the example of the trinary relations ’donation’ and ’debt’. If we look at the early and middle Platonic dialogues we find that relation plays a much larger and sophisticated role than in does in Aristotle’s Organon. The logic of relations does not appear to be very worked at our and was perhaps consciously and deliberately neglected. Indeed there are some natural questions that could be asked (both now and in antiquity). For example: how do definitions work for relations ? What are the differences and how are they joined to the respective proximate genera ? The search for the definition of relatives clearly figures in the Platonic dialogues, for example in the Lysis the definition of friendship is sought after. A definition of a relation can necessarily involve sentences of arbitrary complexity involving other relations which do not fall under the general scheme for monadic predicate terms of book VI of the Topics (at least for its most simple familiar cases). Indeed traces of such unorthodox definitions are to be founded scattered throughout this work222222for example the definition of Justice as Courage and Temperance (150a-151b).. Even the most basic relations such as that of ’being a paternal grandfather’ do not seem to be organized neatly into a proximate genus $+$ difference format. In the Lysis the binary relations of friendship $\Phi$ and love $L$ are considered. The following hypothetical definition is put forward:
@@
The Lysis (see appendix 2) deploys relatively complex statements and reasoning about relations - and none of the proposed definitions seem to fit the genus $+$ difference template. We can wonder why Aristotle did not take this more into consideration in the Topics and Analytics. However there are certain interesting passages which shed light on Aristotle’s view of relations. Consider the passage: 145a14-15: tôn gar pros ti kai hai diaphorai pros ti. Aristotle claims that differences of the relation ’science’ such as ’practical’, ’theoretical’ and ’productive’ are themselves relations. It is not immediately clear how to express this in ASOL (but see the discussion at the end of this section). The discussion in 142a is interesting for it appears to concede that the rule of defining the less know by the better known may fail for relatives such as ’half’ and ’double’. Such relatives must be defined in terms of their correlatives, the double in terms of the half (to diplasion hêmiseos, 142a26).
@@
In appendix 1 we list some of the most important ’logical’ topics in Book IV including (but not limited to) those discussed in this section. We now investigate the structure of the noun phrases used as definitions

### v17

@@ 4.1 Basic framework
which is topic 5 of Chapter 1 listed in appendix 1.
Although the topics are usually listed in an ad hoc fashion it is clear that they are organized into closely knit units bound by logical equivalence or implication. ESOL s required both to formulate the topics and to explain how Aristotle might have established the inferences and equivalence therein. We might call the working out of these implicit inferences the ’metalogic’ of the Topics. Our objective here is not to undertake an exhaustive formal interpretation of all the logical topics in ESOL. Rather it is to go through a selection of key topics mainly from book IV but also from books V and VI which provide evidence for the (extended) second-order nature of the metalogic. For convenience instead of referring directly to the corresponding passage in the text of the Topics for each topic discussed we make use of the lists compiled by Owen in [15]: a formalized version of a portion of them is found in Appendix 1. Let us start with book IV and $\gamma$. We find therein topics which state directly or are stated to be inferred from axioms which express the fact of $\gamma$ having transitive, symmetric and anti-reflexive properties:
@@
as can be verified in appendix 1. The first topic is clearly an ancestor of the Barbara syllogism of the Analytics. Perhaps $\mu$ descends from $\delta$ since the same difference can operate on distinct genera so that $\delta AB$, $A$ is a difference of genus $B$ means that ’some $B$ are $A$’ (a topic discussed further ahead has certainly some connection to the syllogism Darii). There is the important tree-like property:
@@
which is the first topic in chapter 2 of book IV. Aristotle refers to $\gamma AB$ also as ’$A$ participating in $B$’ or ’$A$ receiving the definition of $B$’. Aristotle mentions important axioms in the context of the derivation of refutation or establishing topics. This gives us insight into the metalogic of the topics. For instance in 122b Aristotle infers a topic using the tree-like property and the topic can itself be represented as an inference in sequent form:
@@
Aristotle has the concept of proximate genus which can be easily expressed in ESOL
@@
and so too can the important postulate that every genus must have more than one proximate species (a species of which it is the proximate genus):
@@
But the problem here is to account for the ’equality’ relation which evidently is not the same as extensional equality $_{e}$ but is stronger:
@@
but we do not have (as we argued previously) the reciprocal implication. We could tentatively posit that we have a second-order version of Leibniz’s law for ==(see the topics in 152a which suggests particular cases of this law). In particular in book VII 152a31-33 we read (tr. W. A. Pickard):
@@
Aristotle appears to be stating that $A=B\rightarrow(A=C\rightarrow B=C)$ since
$\sim(A=C\rightarrow B=C)\rightarrow\sim A=B$
Axioms ($\Gamma$) and (I) are stated as a topics. So too are often the counter-positives of some of the topics considered above
All the relations above are exclusive of each other. For instance we have the topic $\sigma AB\rightarrow\sim\gamma AB$ - but this can be derived from ($\Sigma$) and ($\Gamma$). We also have that $\sim\Sigma AA$. There is a topic which seems to state that if $\gamma AB$ then $B$ cannot be the difference of a genus of $A$:
@@
If $\gamma AB$ then there must be some difference $C$ of $B$, $\delta CB$ which ’is said’ of $A$. Aristotle has a kind of unspecific general kind of ’predication’ wherein it is not obvious whether it is supposed to fall under one of the predicabilia or not. We take this unspecified predication to be just extensional containment $A\subset B\leftrightarrow\forall x(Ax\rightarrow Bx)$. We might express this previous topic as
@@
From the above we derive the refutation topic:
@@
We will not analyze here the question of supreme genera and infima species but the order-theoretic properties involving $\gamma$ could be easily expressed in ESOL though other implied properties such as the finiteness of ascending chains of subordinate genera would have to make use of an extension of ESOL.
We have not addressed the ontological categories which feature in the Topics. The order on terms induced by $\gamma$ can in general have more than one maximal element. And we can ask if it is correct to interpret first-order individuals as individual substances. There is the problem that Aristotle seems to distinguish different kinds of situation for which $Ax$, for which a predicate term $A$ holds of an individual substance $x$, for example, if $A$ is predicated according to essence or as an accident. This would suggest that we need an extra predicate term symbol with second-order and first-order arguments $\gamma^{i}Ax$ (for which $\gamma^{i}Ax\rightarrow Ax)$ to make such ontological predication distinctions . A general approach in ESOL to the theory of predication in the The Categories might be something as follows (we will not discuss here the question of whether some of the other categories besides relation are in fact also relations: knowledge is a state of a subject but also, it seems, a relation). We postulate nine different categories of predication $\mathcal{P}^{i}Tx$ where we put $\mathcal{P}^{1}Tx\leftrightarrow Tx$ as corresponding to ousia while the others correspond to the other Aristotelian categories. The tenth will correspond to relation and must split into signatures of the form $\mathcal{P}^{10}R^{2}xy$ and $\mathcal{P}^{\prime 10}R^{2}xT$. For consider ’knowledge’ and ’knowledge of colours’. Knowledge corresponds to the state of a subject but also to a relation, and this relation can regard not one or many primary substances but other categories: if a subject $a$ knows the colour ’blue’ then this is not the same as knowing all blue substances. For another perspective on the categories see the additional notes in the final subsection of this section.
A central role in the Topics is played be a kind of ’semantic’ term negation operator (opposition operator), operating within the proximate species of a given genus. This must be a second-order function applied (primarily) to monadic predicate terms. Its properties are
@@
For instance the opposite of the color white is black. A proximate species of a genus may not have an opposite; this situation is expressed as $A^{\bullet}=A$. Noteworthy topics are
@@
In order to formalize the theory of definition we cannot do without a binary function on monadic predicate terms which corresponds grammatically to a modifier in a noun-phrase, for instance an adjective attached to a head noun. We study in more detail the grammatical aspect of definition terms in the next subsection. We denote such a function by means of concatenation and impose the axiom which is implicitly assumed by Aristotle:
@@
Since we do not want semantically meaningless constructs like ’green numbers’ it makes sense to define a further second-order monadic predicate allowing us to express non non-empty extension (cf. also (NEmp) in section 3):
@@
From this and the axiom for our concatenation operation (Concat) we can deduce the that immediate species of a genus have disjoint extensions:
@@
It is important to note that equality here is not to be read syntactically but semantically. Thus $D=BC$ is to be read as ’the concept $D$ has as parts concepts $B$ and $C$’, the paradigm being ’rational animal’. It is a fundamental problem whether a term can be said to be equal to its definition (with this notion of equality) and whether Leibniz’s law holds for this notion of equality. It could be that two distinct kinds of equality are required. It is true however that definition though the substitution of a term by its definition features in several topics141414152a-152b is very relevant for term equality. Aristotle announces topics which express the transitivity of == as well as topics recalling Leibniz’s law. In 153a he seems to say that there is equality between a term and its definition but this is in itself not enough to show that it is a definition. The definiens must satisfy further conditions. The passage 140b32-141a1 is also interesting. Aristotle states that a definition is identical to the term it defines: esti de tauton tô(i) anthrôpô(i) zô(i)on pezon dipoun. However in 143a15-28 where Aristotle appears to admit that either giving the immediate genus and the difference or a non-immediate genus and all corresponding differences (the ’intension’) as both valid forms of definition which leads us to inquire exactly what the nature of the purported ’uniqueness’ of the definition. Does this means that both these forms are to be regarded as the same definition ?. From this last definition we can evidently derive (in ESOL) numerous refutation topics - and these are indeed topics in book VI. As Aristotle says: it is much easier to refute a definition than to establish one.
A standard view would be that if $hA(DC)$ then $C$ must be uniquely determined (this in fact follows from the above axioms). There is however a problem suggested by the passage 148b dealing with ’complex terms’151515this notion of ’complex term’ appears to be primarily grammatical and not logical or semantic as is our concatenation operation, however the precise relation between the two remains illusive as does the interpretation of the passage in question. Are we to understand the crucial non-circularity property of the definiens in semantic or syntactic terms ? But semantically does it even make sense to have a situation in which $A=AB$ ? .
@@
It is interesting how definitions relates to the opposition operator. If we have a definition $DG$ of $A$ what can we say in general about the definition of $A^{\bullet}$ ? In 153b Aristotle says that there are three cases to consider for the definition for $A^{\bullet}$ depending on whether $G^{\bullet}=G$ or $D^{\bullet}=D$. The possibilities are either $D^{\bullet}G^{\bullet}$, $D^{\bullet}G$ or $DG^{\bullet}$. For example justice is virtue of the soul and its opposite injustice is the vice of the soul, soul having no opposite.
4.2 On relations
We now discuss Aristotle’s relation terms concerning which section 4 of book IV of the Topics is of particular interest. A natural approach is to represent relation terms by binary predicate terms in ESOL. But then we need a corresponding version of $\gamma$ for such arguments, there is no way to deal uniformly with the predicabilia when predicate terms of different arities are involved. This is also enforced by the topic in book IV section 1 (120b) which requires that species and genus belong to the same category, although Aristotle admits later in section 4 that there might be exceptions (124b) as for the case of ’grammatical’ (monadic) as a species of the genus ’knowledge’(relation). In our case the application of the aforementioned topic (stated also explicitly in (124b)) is simply that there must be two species-genus relations, one $\gamma$ for monadic predicates and another $\gamma_{2}$ for binary predicate terms. We need to postulate analogous axioms analogous to those for $\gamma$ such as
@@ 4.2 On relations
where $R,S$ and $T$ are binary predicate term variables. The logic of relation terms is quite complex in Aristotle and bound up with Greek syntax in a way which renders interpretation difficult. A relation term $R$ has its ’reciprocal’ or ’inverse’ which we can interpreted by a second-order function $R^{\ast}$. It is a challenge is to interpret the passage in (124b) about the double ($D$) being the double of the half. It is plausible that we are in this case in the presence of relations for which
$\forall x,x^{\prime},y,y^{\prime}(Rxy\&Rxy^{\prime}\rightarrow y=y^{\prime})\&(Rxy\&Rx^{\prime}y\rightarrow x=x^{\prime})$
@@
In the passages that follow Aristotle discusses the necessary agreement between the case governed for terms in the $\gamma_{2}$ relation. For example if $\gamma_{2}RS$ then if $R$ takes the genitive then so does $S$. More interesting from a logical point of view is the passage mentioning the number of complements, this suggests that Aristotle was conscious of the necessity of the general agreement between the arity of the predicate terms in the species-genus relation, so that we should consider a whole sequence of such relations $\gamma_{1}=\gamma,\gamma_{2},...,\gamma_{i},.$... where $\gamma_{n}$ serves for $n$-ary predicate variables. Aristotle uses the example of the trinary relations ’donation’ and ’debt’. If we look at the early and middle Platonic dialogues we find that relation plays a much larger and sophisticated role than in does in Aristotle’s Organon. The logic of relations does not appear to be very worked at our and was perhaps consciously and deliberately neglected. Indeed there are some natural questions that could be asked (both now and in antiquity). For example: how do definitions work for relations ? What are the differences and how are they joined to the respective proximate genera ? The search for the definition of relatives figures in the Platonic dialogues, for example in the Lysis the definition of friendship is sought after. A definition of a relation can necessarily involve sentences of arbitrary complexity involving other relations which do not fall under the general scheme for monadic predicate terms of book VI of the Topics (at least for its most simple familiar cases). Indeed traces of such unorthodox definitions are to be founded scattered throughout this work161616for example the definition of Justice as Courage and Temperance (150a-151b).. Even the most basic relations such as that of ’being a paternal grandfather’ do not seem to be organized neatly into a proximate genus ++ difference format. In the Lysis the binary relations of friendship $\Phi$ and love $L$ are considered. For example the following hypothetical definition is put forward:
@@
The Lysis deploys relatively complex statements and reasoning about relations - and none of the proposed definitions seem to fit the genus ++ difference template. We can wonder why Aristotle did not take this more into consideration in the Topics and Analytics. However there are certain interesting passages which shed light on Aristotle’s view of relations. Consider the passage: 145a14-15: tôn gar pros ti kai hai diaphorai pros ti. Aristotle claims that differences of the relation ’science’ such as ’practical’, ’theoretical’ and ’productive’ are themselves relations. It is not immediately clear how to express this in ESOL (but see the discussion further ahead). The discussion in 142a is interesting for it appears to concede that the rule of defining the less know by the better known may fail for relatives such as ’half’ and ’double’. Such relatives must be defined in terms of their correlatives, the double in terms of the half (to diplasion hêmiseos, 142a26).
@@
As already mentioned, in Appendix 1 we list some of the most important ’logical’ topics in Book IV including (but not limited to) those discussed in this section. We now investigate the structure of the noun phrases used as definitions, which is relevant also to a refinement of our ’concatenation’ operation.
4.3 The grammatical structure of definition terms